OpenStudy (anonymous):

Could somebody help e with this limit ? D: lim 1-cosx/x^2/3 approaches x-> 0 please:D

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{1-\cos(x)}{x^\frac{2}{3}}\] is this right?

OpenStudy (freckles):

if so try multiplying top and bottom by x^(1/3)

OpenStudy (freckles):

you should know a famous trig limit you can use

OpenStudy (anonymous):

dang can't figure out how to typeset l\Hopital not that it is needed...

OpenStudy (anonymous):

yeah is the first one ! But I can't Im new in this topic

OpenStudy (freckles):

do you know what this equals: \[\lim_{x \rightarrow 0}\frac{1-\cos(x)}{x}=?\]

OpenStudy (anonymous):

zero right?

OpenStudy (freckles):

yes

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{1-\cos(x)}{x^\frac{2}{3}} \cdot \frac{x^\frac{1}{3}}{x^\frac{1}{3}}\] can you figure out why I want to multiply top and bottom by x^(1/3)?

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{1-\cos(x)}{x} \cdot x^\frac{1}{3}\]

OpenStudy (freckles):

you can write this as \[\lim_{x \rightarrow 0}\frac{1-\cos(x)}{x} \cdot \lim_{x \rightarrow 0}x^\frac{1}{3}\] since both limits exist

OpenStudy (freckles):

you already told me the first limit in the product equal 0 the second function is continuous at x=0 so you can use direct sub for the second limit in the product

OpenStudy (anonymous):

so the final answer will be 0 because if I multiply \[x^{1/3} \] in 1-cos x , that will be zero and the denominator will be x so when i evaluate it will be 0/ 1=0 right?

OpenStudy (freckles):

well (1-cos(x))/x->0 and so you have 0*0 which is 0

OpenStudy (anonymous):

Thanks finally I understood :D